OpenStudy (anonymous):

solve each equation for solutions in the interval 0≤x<2π

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a) 2sinxcosx=√(3) sinx

OpenStudy (anonymous):

2-√(3)=sinx okay r u agree?

OpenStudy (anonymous):

coz sinxcosx=1

OpenStudy (anonymous):

\[cos(x)=\frac{\sqrt{3}}{2}\] that's what I get after dividing by sin(x) because it is redundant and dividing by two.

OpenStudy (anonymous):

okay why just take sin(x)

OpenStudy (anonymous):

this forms seems much easier to me to solve, you can get rid of the sin(x), since it is represented on the lefthand- and also on the righthand side. Therefore it's just another scalar term for the solution. Remember that \[2\sin(x)\cos(x)=\sin(2x)\]

OpenStudy (anonymous):

yaeh

OpenStudy (anonymous):

What you obtain will be: \[x_1= \frac{\pi}{6} \] You can get the second solution by looking at the unit circle. \[x_2= \pi - x_1 = \frac{5 \pi}{6} \]

OpenStudy (anonymous):

i ended up getting x=π/3, 4π/3, 3π/4, 7π/4

OpenStudy (anonymous):

arm sorry I messed the last term up, it's a full rotation of course, therefore 2 pi

OpenStudy (anonymous):

hmm some of your solution don't seem right to me @charliem07, check them by back substitution.

OpenStudy (anonymous):

I get \[x_1 = \frac{\pi}{6} \\ x_2 = 2\pi - x_1 =\frac{11 \pi}{6} \]

OpenStudy (anonymous):

and then you can sum them up, using the period.

OpenStudy (anonymous):

@Spacelimbus i thought you list them since they are in interval

OpenStudy (anonymous):

ohhhh i see what you are saying thanks

OpenStudy (anonymous):

you are welcome